ABCD is a quadrilateral with G and H the mid points of line DA and DC respectively. Show that line AB + line DC = 2GH


1
Expert's answer
2021-06-04T12:51:31-0400


Given a quadrilateral "ABCD" with "G" and "H" the midpoints of line "DA" and "DC" respectively. 

Consider triangle "ACD"

 A midsegment of a triangle is a line segment which joins the midpoints of two sides of the triangle. It is parallel to the third side and has a length equal to half the length of the third side. Hence "2GH=AC."


Consider the square "ABCD" with side "1"


"AB=BC=CD=DA=1"

Consider right triangle "ACD." By the Pythagorean Theorem


"AC^2=AD^2+DC^2"

"AC=\sqrt{2}"

"2GH=AC=\sqrt{2}"

"AB+DC=1+1=2>\sqrt{2}=2GH"

Hence


"2GH<AB+DC."


The statement  line AB + line DC = 2GH for a quadrilateral is False in our case.